Is the function f:R→R defined by f(x)=cos(2x+1) invertible?Give reasons.Can you modify domain and codomain of f such that f becomes invertible?

Grade:12th Pass

3 Answers

jitender lakhanpal
12 years ago

Dear Menka,

no the function is not invertible as the function f(x) is not BIJECTIVE as it is not one to one

as there are many values of x for which there exist values of -1 to 1 the range of the function

and the function is also not onto as the codomain is all real number R but the range of f(x) is -1 to 1

and for onto the codomain must be equal to range .

as we know the necessary and sufficient condition for the function to be invertible should be 

that the function must be bijective .

to make function invertible the codomain must be equal to [-1 , 1]

and domain must be 

0 <= 2x + 1 <= pie

now by this relation we get

-1/2 <= x <= (pie - 1 )/2

so the domain must be equal to -1/2 <= x <= (pie - 1 )/2

and codomain is  [-1 , 1]. to make function invertible

 

 

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

YOUR NAME 

jitender

Ashwin Muralidharan IIT Madras
12 years ago

Hi Menka,

 

Here's the answer in the scanned copy:

Hope that helps.

 

All the best.

Regards,

Ashwin (IIT Madras).

jitender lakhanpal
12 years ago

hi ashwin 

i did not get the first part of the solution was it asked 

and i think i have done the problem in simpler way